HETAV PATEL

Grade 11,

find four members inGP whose sum is 85 and product is 4096

find four members inGP whose sum is 85 and product is 4096

Grade:9

2 Answers

Arun
6 years ago
Dear Priya
 
Let 4 numbers in GP are = a/r³, a/r, ar, ar³
Now
a(1/r³ + 1/r + r + r³) = 85 
And 
a⁴ = 4096
a = 8
Now you can find r,
After that also youncan find 4 numbers.
 
Regards
Arun (askIITians forum expert)
 
sahil
6 years ago
let the term be x/r^3 ,x/r ,xr and xr^3now product =4096 =x/r^3×x/r×xr+xr^3=x^4 so x=4096^1/4=8now terms are 8/r^3+8/r+8r+8r^3sum =85 solve to get r=2so terms are 8/2^3,8/2,8×2,8×2^3 lastly terms are 1,4,16,64If correct please give approval